OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y =1/4x + 7 and that pass through the point (-8, 9). a) parallel: y =1/4x - 11 perpendicular: y = -4x + 23 b) parallel: y = -4x + 11 perpendicular: y =1/4 x - 4 c)parallel: y = 4x - 7 perpendicular: y =1/4x - 7 d)parallel: y =1/4x + 11 perpendicular: y = -4x - 23

OpenStudy (anonymous):

help?

OpenStudy (aroub):

Oh my x_x This what I'm going to do, I'll explain to you the first one and you do the rest and I'll check them, deal? :)

OpenStudy (anonymous):

okay deal :)

OpenStudy (aroub):

Do you know what does parallel and perpendicular mean here?

OpenStudy (anonymous):

a little bit not so much :/

OpenStudy (aroub):

It's okay, what do you know?

OpenStudy (anonymous):

okay i know that perpendicular lines intersect at a 90 degree angle and parallel lines dont intersect

OpenStudy (aroub):

Okay, the meaning in here for both is all about the slope. Parallel means that they have the same slope while in perpendicular the slope is the opposite reciprocal of the original slope.

OpenStudy (anonymous):

ohh okay i got it.

OpenStudy (aroub):

For a) parallel: y =1/4x - 11 ; (-8,9) slope= 1/4 ; x=-8 , y=9 ( I got x and y from (-8,9) ) To solve such problem you will have to use the slope intercept form, which is: y=mx+b, where m is the slope. Now all you got to do is substitute to find b. => 9=1/4(-8)+b 9=-2+b 11=b since we found" b" we can now find the new equation which is parallel to the original equation. How? By using the slope intercept form, again! y=mx+b, you always leave the x and y! Just substitute for m and b. => y=1/4x+11 Got it?

OpenStudy (anonymous):

yes thank you so much :)

OpenStudy (aroub):

Can you do the perpendicular now? They're the exact same steps but with different slopes of course. Remember the slope of the perpendicular is the opposite reciprocal.

OpenStudy (aroub):

I say -4, agree?

OpenStudy (anonymous):

yes now i can do it and i agree with -4

OpenStudy (aroub):

Great ^_^